An investor purchases a stock for $28 and a put on the stock for $0.40 with a strike price of $24. She also sells a call on the same underlying stock for $0.40 with a strike price of $30 and with the same expiration date. What is the value of her portfolio, net of the proceeds from the options, if the stock price ends up at $35 on the expiration date?
1
Expert's answer
2020-07-28T10:49:35-0400

The collar involves purchasing a put for 0.40 and selling a call far 0.40. The initial outlay is 0.

St=35

Value at expiration = Value of call + Value of put + Value of stock=0+(28-35)+0+35=28

Final value:

28-0=28